OpenStudy (anonymous):

If radius of circle is 12cm, find the length of arc ABC (drawing in comments)

OpenStudy (anonymous):



OpenStudy (anonymous):

\[2\pi \times radius \approx circumference; 24\pi \approx circumference. So, 24\pi \times 80/360 = \] length of arc, right?

OpenStudy (ranga):

That would be the length of the arc AC. Here they are asking for arc ABC (the longer arc). So it should be 2(pi)r minus your earlier answer: 24(pi) - 24(pi) * 80/360 = 24(pi)[1 - 80/360] = 24*pi*280/360 = 56(pi)/3